# Structured Episodic Event Memory

### v1

@@ 4.4 Implementation Details
We standardize the backbone models across all methods to ensure a fair comparison. We primarily employ Qwen3-Next-80B-A3B-Instruct Yang et al. (2025) for both information extraction and downstream question answering tasks. To further demonstrate the model-agnostic robustness of the SEEM framework, we further conduct additional experiments using GPT-OSS-120B Agarwal et al. (2025) as the backbone. Detailed results for this cross-model validation are provided in Appendix A.1. Regarding retrieval configurations, we align the hyperparameters based on the granularity of the retrieved units. For standard RAG baselines that rely on dense retrieval, we set the retrieval count $k$ to 5, fetching the top-5 original interaction messages. Similarly, for the memory-augmented baselines Mem0 and A-MEM, we retrieve the top-10 processed memory chunks. For HippoRAG 2, which operates on a session-based retrieval logic, we utilize the top-5 retrieved chunks to construct the context for final response generation. In our proposed SEEM framework, we configure the system to retrieve the top-5 relevant text chunks alongside their associated episodic memories to construct the reasoning context. To balance narrative continuity with information density, we employ a selective RPE strategy. Specifically, the total size of the final expanded evidence set $\mathcal{P}_{final}$ is restricted to at most twice the initial retrieval budget.

### v2

@@ 4.1 Datasets
Method LoCoMo LongMemEval
F1 EM $J$ F1 EM Acc.
Context (Relevant Only) 38.60 11.40 69.28 22.60 3.20 68.60
NV-Embed-v2 (Lee et al., 2025) 46.10 29.10 51.56 16.53 2.00 57.20
Memory-based Frameworks
HippoRAG 2 (Gutiérrez et al., 2025) 44.60 27.90 50.96 34.82 25.10 46.59
SEEM (Ours) 48.60 30.80 56.11 45.50 29.60 60.80
Table 3: Performance comparison on LoCoMo and LongMemEval using the Pangu-Embedded-7B backbone. The best results are highlighted in bold.
@@ 4.4 Implementation Details
We standardize the backbone models across all methods to ensure a fair comparison. We primarily employ Qwen3-Next-80B-A3B-Instruct Yang et al. (2025) for both information extraction and downstream question answering tasks.To further validate the model-agnostic robustness and efficiency of the SEEM framework, we additionally incorporate Pangu-Embedded-7B Chen et al. (2025), a compact model with 7 billion parameters, as a secondary backbone. This allows us to observe whether the episodic memory benefits scale down effectively to smaller-parameter models. Furthermore, cross-model validation using GPT-OSS-120B Agarwal et al. (2025) is provided in Appendix A.1. Regarding retrieval configurations, we align the hyperparameters based on the granularity of the retrieved units. For standard RAG baselines that rely on dense retrieval, we set the retrieval count $k$ to 5, fetching the top-5 original interaction messages. Similarly, for the memory-augmented baselines Mem0 and A-MEM, we retrieve the top-10 processed memory chunks. For HippoRAG 2, which operates on a session-based retrieval logic, we utilize the top-5 retrieved chunks to construct the context for final response generation. In our proposed SEEM framework, we configure the system to retrieve the top-5 relevant text chunks alongside their associated episodic memories to construct the reasoning context. To balance narrative continuity with information density, we employ a selective RPE strategy. Specifically, the total size of the final expanded evidence set $\mathcal{P}_{final}$ is restricted to at most twice the initial retrieval budget.
@@ 5.1 Main Results
Efficacy on Parameter-Constrained Backbones.
The results in Table 3 reveal a compelling trend: SEEM effectively narrows the performance gap between small-scale models and their larger counterparts. Even when powered by Pangu-Embedded-7B, SEEM consistently outperforms all baseline frameworks, including those utilizing more complex retrieval-augmented strategies. Specifically, the significant lead in the reasoning assessment ($J$) on LoCoMo suggests that SEEM’s episodic memory provides higher information density and narrative coherence, which is crucial for smaller LLMs that often struggle with "distraction" in long-context windows. Furthermore, the robust accuracy on LongMemEval confirms that SEEM does not rely on massive parameter counts to maintain long-term context; instead, it provides a specialized memory scaffolding that enables a 7B model to achieve reasoning depth previously reserved for much larger architectures. This highlights SEEM as a model-agnostic, parameter-efficient solution for long-range agentic memory.